OpenStudy (anonymous):

A red ribbon is tied tightly around the earth at the equator. How much more ribbon would you need if you raised the ribbon 1 ft above the equator everywhere? (You don't need to know the radius of the earth to solve this problem.) ~Could you also include an explanation of how you went about solving this. Thanks!

OpenStudy (anonymous):

I can give you an answer as a factor of the current lenght of the ribbon.

OpenStudy (anonymous):

Okay~ that may help

OpenStudy (anonymous):

let original radius of earth at equator = R ft ribbon required = circumference = C1 = 2πR when it is lifted 1 ft above the equator, now radius = (R + 1) ft now ribbon required = circumference = C2 = 2π(R + 1) = 2πR + 2π Increase in length of ribbon = C2 - C1 = (2πR + 2π) - 2πR = 2π = 2*3.14 = 6.28 ft

OpenStudy (anonymous):

@ noru pls see my solution....... no values are required........

OpenStudy (anonymous):

i saw... i confused my formula you are right

OpenStudy (anonymous):

Yeah~ Harkirat that makes sense. Thank you so much! I appreciate it! :)

OpenStudy (anonymous):

u r always welcome....☺